OpenStudy (anonymous):

Express each quotient in the simplest form. t+4 / t-4 divide t+4 / t-4 A. 1 / 16 B. 1/4 C. 1 D. 2(t+4) / t - 4 help please. will give a medal :)

OpenStudy (mathmate):

hint: \(\frac{a}{b}\div \frac{c}{d} = \frac{ad}{bc}\)

OpenStudy (anonymous):

okay i feel its A i just have a feeling since 4x4 is 16 am i wrong?

OpenStudy (i0iz0gangster):

well remember the last problem you did :) try to solve the equation and lets see your answer

OpenStudy (anonymous):

idk pellet >.<

OpenStudy (i0iz0gangster):